Q1:one year ago pandit was three times his sisters age. next year he will be only twice her age. how old will be pandit after5 years.


Q2:a man was arrested for exeeding the speed limit by 10 miles an hour. a second man was charged with exceeding the same limit by twice as much. the second man was driving 35 miles per hour then what was the speed limit.


Q3:


A princess is as old as the prince will be when the princess is twice the age that the prince was when the princess’s age was half the sum of their present ages.


 


What are their ages?

Grade:12

1 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear naidu,

Let age of pandit be x years at present

and age of his sister be y years at present

from the question

x-1 = 3(y-1)

x+1 = 2(y+1)

solving these two equations

you get y = 3 and x = 7

after 5 years , age of pandit will be = 12 yrs

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB